Facts: You purchased a piece of equipment on January 1, 2019 for $53,000. It has a $3,000 salvage value and a 5 year useful life. It’s life in units of output is 100,000 hours. Prepare the depreciation schedules for the entire useful life (e.g. You will show all the years for the asset) as follows:

  1. Straight line
  2. Sum of the year’s digits
  3. Units of production (use the following use: yr1: 18,000; yr2: 22,000; yr3: 28,000 yr4: 20,000; yr5: 12,000
  4. 200% Declining balance
  5. Straight line if on January 1, 2021, you decide you actually have 4 more years of useful life.
Answer #1

Straight line depreciation

Cost= 53,000 salvage value =3,000

Life = 5 years,

Dep per year= (Cost - salvage value)/No of years

= (53,000 - 3,000)/5 years

= 10,000

Depreciation schedule

Year Book value year start Depreciation Expenses Accumulated depreciation Book Value end
2019 53,000 10,000 10,000 43,000
2020 43,000 10,000 20,000 33,000
2021 33,000 10,000 30,000 23,000
2022 23,000 10,000 40,000 13,000
2023 13,000 10,000 50,000 3,000

Sum of year digit method

Depreciation calculated as on year 1 =( cost-salvage value)× remaining life /total of years of life

Total of life = 5+4+3+2+1= 15

Dep of year 1 = (53,000- 3,000)×5/15

= 16,667

Year 2 = (53,000-3,000)×4/15

= 13,333 and so on

Depreciation schedule

Year Book value year start total cost Depreciable Depreciation percentage depreciation expense Accumulated depreciation Book value year end
2019 53,000 50,000 33% 16,667 16,667 36,333
2020 36,000 50,000 27% 13,333 30,000 23,000
2021 23,000 50,000 20% 10,000 40,000 13,000
2022 13,000 50,000 13% 6,667 46,667 6,333
2023 6,333 50,000 7% 3,333 50,000 3,000

Unit of production Method

Dep = (cost-salvage value)× no of output used in year/total output

Year 1 = (53,000- 3,000)× 18,000/100,000=9,000

Year 2= (53,000-3000)×22,000/100,000=11,000

Year 3= (53,000-3,000)×28,000/100,000=14,000

Year4= (53,000-3,000)×20,000/100,000=10,000

Year 5=(53,000-3,000)×12,000/100,000=6,000

Depreciation schedule

Year book value year start Depreciation Expenses Accumulated depreciation book value ending
2019 53,000 9,000 9,000 44,000
2020 44,0000 11,000 20,000 33,000
2021 33,000 14,000 34,000 19,000
2022 19,000 10,000 44,000 9,000
2023 9,000 6,000 50,000 3,000

200% declining balance

Dep rate = cost/useful life

= 53,000/5= 10,600

Rate = 10,600×100/53,000

=20%

DDB rate = 20% × 2=40%

Depreciation schedule

Year book value year start

Depreciation percent

Depreciation expense Accumulated depreciation book value year end
2019 53,000 40% 21,200 21,200 31,800
2020 31,800 40% 12,720 33,920 19,080
2021 19,080 40% 7,632 41,552 11,448
2022 11,448 40% 4,579 46,131 6,889
2023 6,869 40% 2,748 48,879 4,121

Book value on Jan 1 2021 as per straight line=33,000

Remaining useful life= 4 years

Dep per year= (33,000- 3,000)/4 years

=7,500

Depreciation schedule

Year Book value year start Depreciation Expenses Accumulated depreciation book value year end
2019 53,000 10,000 10,000 43,000
2020 43,000 10,000 20,000 33,000
2021 33,000 7,500 27,500 25,500
2022 25,500 7,500 35,000 18,000
2023 18,000 7,500 42,500 10,500
2024 10,500 7,500 50,000 3,000